OpenStudy (anonymous):

Find the smallest value of n so that 180n is a cube number (show working out please)

OpenStudy (anonymous):

n = 150

OpenStudy (anonymous):

180n = 3^2 * 2^2 * 5 * n so n must be equal to 3 * 2 * 5 ^2

OpenStudy (anonymous):

so do we just write the ans as 150 or do we need to times with it with 180 (180x150) or 150 is the ans? thanks for replying as i have an exam tmr

OpenStudy (anonymous):

yes the answer is 150.

OpenStudy (anonymous):

if you dont mind can you please help me with this question ( Find the smallest value of n such that 1800n is a perfect cube) srry as i really suck in math :3

OpenStudy (anonymous):

1800n = 3^2 * 2^3 * 5^2 * n smallest value for n = 3 * 5 = 15

OpenStudy (anonymous):

Thank you very much!:)

OpenStudy (anonymous):

good luck with your exam :)